OpenStudy (anonymous):

What is the percent for 11/50

OpenStudy (anonymous):

11/50x100

OpenStudy (anonymous):

u do that for all fractions

OpenStudy (anonymous):

therefore 22%

OpenStudy (anonymous):

wat???

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

not really its : 11/50 * X = Y/100 for this case its *2 11/50*2 = 22/100 22%